# MatFormer: Nested Transformer for Elastic Inference

### v1

@@ Abstract
Transformer models are deployed in a wide range of settings, from multi-accelerator clusters to standalone mobile phones. The diverse inference constraints in these scenarios necessitate practitioners to train foundation models such as PaLM 2, Llama, & ViTs as a series of models of varying sizes. Due to significant training costs, only a select few model sizes are trained and supported, limiting more fine-grained control over relevant tradeoffs, including latency, cost, and accuracy. This work introduces MatFormer [^1] MatFormer stands for [Uncaptioned image] Matryoshka Transformer due to the model’s inherent nested nature., a nested Transformer architecture designed to offer elasticity in a variety of deployment constraints. Each Feed Forward Network (FFN) block of a MatFormer model is jointly optimized with a few nested smaller FFN blocks. This training procedure allows for the Mix’n’Match of model granularities across layers – i.e., a trained universal MatFormer model enables extraction of hundreds of accurate smaller models, which were never explicitly optimized. We empirically demonstrate MatFormer’s effectiveness across different model classes (decoders & encoders), modalities (language & vision), and scales (up to 2.6B parameters). We find that a 2.6B decoder-only MatFormer language model (MatLM) allows us to extract smaller models spanning from 1.5B to 2.6B, each exhibiting comparable validation loss and one-shot downstream evaluations to their independently trained counterparts. Furthermore, we observe that smaller encoders extracted from a universal MatFormer-based ViT (MatViT) encoder preserve the metric-space structure for adaptive large-scale retrieval. Finally, we showcase that speculative decoding with the accurate and consistent submodels extracted from MatFormer can further reduce inference latency.
@@ 1 Introduction
Figure 1: MatFormer introduces nested structure into the Transformer’s FFN block & jointly trains all the submodels, enabling free extraction of hundreds of accurate submodels for elastic inference.
@@
Large Foundation models (Anil et al., 2023; OpenAI, 2023; Dehghani et al., 2023) are deployed in a variety of settings like real-time response on mobile phones or in batch setting on multi-cluster GPUs for web-scale serving. To handle such varied settings, each model family provides a few independently trained models of different sizes. In order to cover a wide range of applications, typically these models’ sizes are nearly linear on log-scale. For example, Llama family provides models with 7B, 13B, 33B and 65B parameters (Touvron et al., 2023a).
Such an approach has two key drawbacks: (a) as the models are independently trained, they incur significant overhead for colocation during inference and are not behaviorally consistent with each other which are detrimental to inference optimization techniques like speculative decoding (Leviathan et al., 2023) and model cascades (Wang et al., 2020b), and (b) due to training overhead, practitioners typically train only a few models which do not cover the entire set of downstream use-cases. For example, a deployment setup might, say, have the latency budget to support 40B parameter Llama model, but can only host a 33B variant because the next bigger model (65B) has significantly higher latency. So, one would need to settle for a less accurate model despite the larger latency budget. While model compression approaches aim to address this issue, they typically require additional training for each model that needs to be extracted. Furthermore, when applied to LLMs, these techniques are known to significantly drop the accuracy (Jaiswal et al., 2023).
@@
In this paper, we propose MatFormer, a natively elastic Transformer (Vaswani et al., 2023) architecture that allows for training one universal model which can be used to extract hundreds of smaller submodels without any additional training (Figure 1). MatFormer is a general architecture that can be applied to both encoders and decoders, is domain agnostic, and is compatible with most design choices and training pipelines of large Transformer-based models – LLMs & ViTs.
MatFormer follows the principle of matryoshka representation learning (Kusupati et al., 2022), to introduce nested substructure inside the standard Transformer block. Formally, MatFormer defines a Transformer blocks $T_{i}$, such that, $T_{1}\subset T_{2}\subset\dots\subset T_{g}$, where $g$ is the number of nested transformer blocks, and $T_{i}\subset T_{i+1}$ relation indicates that the parameters of $T_{i}$ are contained in those of $T_{i+1}$.
@@
MatFormer can induce such sub-structure in both the attention and the feedforward network (FFN) blocks of the Transformer (see Figure 1). Consider a FFN block that has $d_{\text{ff}}$ neurons in the hidden layer. Then, MatFormer induces matryoshka structure on these neurons, where $T_{i}$ contains the first $m_{i}$ neurons and $1\leq m_{1}\leq m_{2}\dots\leq m_{g}=d_{\text{ff}}$ represent the number of neurons for each granularity or sub-model. Intuitively, this implies that the first $m_{1}$ neurons are “most significant” neurons as they belong to all the blocks followed by the next $m_{2}-m_{1}$, and so on. We can form a similar sub-structure on the attention heads, with the heads being organized from “most” to “least” significant, where the more significant heads are shared by more sub-models. In fact, we can also introduce this sub-structure in the token embedding ($d_{\text{model}}$) supplied to each Transformer block.
However, in most LLMs and ViTs, the FFN block in the Transformer accounts for more than $60\%$ non-embedding parameters and is responsible for the largest chunk of latency during inference. So, in this work, we focus on inducing the MatFormer’s nested sub-structure in the FFN block. We then stack the individual blocks (for $l$ layers) to form $g$ nested models ($\mathcal{M}_{1\dots g})$ with shared parameters i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{i+1}$. Finally, we jointly train these $g$ models by combining each model’s loss.
@@
This leads to a natural question: can one extract more than $g$ models after inducing the MatFormer structure? Yes, in fact, it is possible to extract exponentially many models. Using the trained MatFormer blocks $T_{1},\dots,T_{g}$ at each layer, one can form new models by Mix’n’Match, i.e., by taking an arbitrary combination of these blocks across layers. For example, in the first layer, one can select $T_{g}$, the largest block, choose $T_{2}$ in the second layer, and so on, forming $g^{l}$ different models. As we explicitly optimized only for $g$ models, instead of the exponentially many models, are the extracted models accurate? Surprisingly, in multiple settings, and for a various model sizes, we observe that the extracted models indeed are accurate, with accuracy scaling with the size of the extracted model.
@@
We train Matformer-based decoder-only Language Models (MatLM) up to 2.6B parameters and observe that: (a) MatLMs explicitly trained with $g$ exponentially spaced granularities (4 in this work) almost match validation loss and one-shot downstream evals of respective $g$ baseline models trained independently from scratch, (b) our extracted models using Mix’n’Match lie on the accuracy-vs-parameters trade-off curve generated by the $g$ explicitly trained models, (c) through multiple scaling experiments we observe that the loss vs compute law for different MatFormer models remains similar to vanilla Transformer models across different granularities and (d) the submodels extracted from MatLM natively have highly consistent behavior that is highly desirable for inference optimizations and deployment across scales.
@@
We further studied MatFormer-based ViT models (MatViT) and have similar observations as MatLM. For example, MatViT-L/16 improves the accuracy of the standard ViT-L/16 model on ImageNet-1K, and the extracted sub-models all match or even perform better than the independently trained baselines. Furthermore, we demonstrate that, due to high consistency, MatViT models can be used as “elastic encoders” for adaptive image retrieval. That is, the metric-space of an image encoded by the universal (i.e. the largest) MatViT model is roughly preserved by the nested submodels. Hence, based on query complexity, system load, and various other considerations, we can use one of the extracted MatViT encoders at inference time for retrieval on a fixed corpus encoded by the universal model – providing over $40\%$ lesser compute overhead with $<0.5\%$ drop in accuracy.
@@
    We introduce MatFormer, which incorporates a nested sub-structure within the standard Transformer and jointly optimizes all the $g$ granularities to produce a single, universal elastic model.
@@
    Employing Mix’n’Match of granularities across layers in a universal MatFormer model yields hundreds of accurate and consistent submodels without any additional training cost (Section 3).
@@ 2 Related Work
A standard Transformer (Vaswani et al., 2023) has become the unifying model architecture for foundation models (Bommasani et al., 2021) across modalities like language (Brown et al., 2020), vision (Dehghani et al., 2023) and audio (Radford et al., 2023). While extremely powerful, the standard Transformer block is not natively elastic in a way that enables large-scale adaptive and flexible deployment across various resource constraints. To cater to the plethora of deployment requirements, existing solutions include training a family of models of varying sizes (Anil et al., 2023; Touvron et al., 2023b), post-hoc efficiency techniques like quantization (Dettmers & Zettlemoyer, 2023), pruning (Lagunas et al., 2021), distillation (Sanh et al., 2019) and mixture of varying capacity experts (MoE) (Zhang & Ma, 2012). However, these solutions often are specific to the single constraint at hand, and require additional training or trade-off memory/compute during inference making them far from being a truly elastic solution for adaptive deployment. Lastly, Transformer based LLMs are often sped-up during inference with techniques like speculative decoding (Leviathan et al., 2023; Chen et al., 2023) – that benefits from the smaller draft & the larger verifier models having similar behavior – or early exiting (Schuster et al., 2022) to enable real-time deployment.
@@
Obtaining multiple smaller models from a single model has been explored in the past (Yu et al., 2018; Yu & Huang, 2019; Cai et al., 2019; Grimaldi et al., 2022) with most works focusing on CNN encoders. For example, OFA (Cai et al., 2019) creates a universal CNN model which is used to extract and finetune submodels for a handful of deployment constraints while slimmable networks (Yu et al., 2018) optimize for limited preset widths and require explicit training to interpolate for a few more intermediate widths (Yu & Huang, 2019). These techniques fall short of being truly elastic and come with significant training overheads. More recently some of them have been extended to Transformer encoders (Chavan et al., 2022; Hou et al., 2020; Salehi et al., 2023) for extracting sub-models in both static or dynamic settings but fail at extending further to decoder-only language models. While not in the weight space, matryoshka representation learning (Kusupati et al., 2022) & FlexiViT (Beyer et al., 2023) showcase elasticity in output & input spaces respectively by smoothly spanning deployment constraints with minimal overhead. MatFormer is inspired by both these lines of work to enable truly elastic and adaptive Transformer-based (decoder & encoder) models that span all the accuracy-vs-compute tradeoff (statically or dynamically) with minimal changes and training overhead (Figure 1). Finalyl, we also point the readers to SortedNet (Valipour et al., 2023), a concurrent work with similar goals applied to encoders, which optimizes many sampled submodels (akin to prior works) unlike MatFormer’s joint optimization of a few (typically 4) nested submodels.
@@ 3.1 MatFormer Structure
MatFormer defines $g$ Transformer blocks $T_{i}$, such that, $T_{1}\subset T_{2}\subset\dots\subset T_{g}$ where $T_{i}\subset T_{i+1}$ indicates that the parameters of $T_{i}$ are contained in those of $T_{i+1}$. While it is possible to impose such a structure on any part of the Transformer, we select the FFN block to define our method and present our experiments, as the model size and computational cost of a Transformer is dominated (around $60\%$ for LLMs and ViTs) by the FFN block (see Appendix B).
The FFN block in Transformer has a single hidden layer with $d_{\text{ff}}$ neurons and both input and outputs in $\mathbb{R}^{d_{\text{model}}}$, and a fixed FFN ratio $:=d_{\text{ff}}/d_{\text{model}}$ (typically $\geq 4$). MatFormer introduces the matryoshka nested structure with $g$ granularities on the hidden representation $d_{\text{ff}}$ of the FFN block. Concretely, a nested sub-block of the Transformer, $T_{i}$ contains the first $m_{i}$ neurons of the FFN and $1\leq m_{1}\leq m_{2}\dots\leq m_{g}=d_{\text{ff}}$ represent the number of neurons for each granularity or sub-model. So depending on the chosen granularity the FFN operation of $T_{i}$ i.e., $T_{i}^{\text{FFN}}$ on an input $x\in\mathbb{R}^{d_{\text{model}}}$ is:
@@
where the weight matrices of FFN are $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{d_{\text{ff}}\times d_{\text{model}}}$ and bias terms are omitted for simplicity. $\mathbf{W}_{1}[0:k]$ denotes the submatrix with the first $k$ rows of $\mathbf{W}_{1}$. Finally, $\sigma$ is a non-linearity often set to GELU (Hendrycks & Gimpel, 2016) or squared ReLU (So et al., 2021). In this work, we chose the $g=4$ exponentially spaced granularities with FFN ratios of $\{0.5,1,2,4\}$ i.e., the nested hidden neurons are of the sizes $\{\frac{d_{ff}}{8},\frac{d_{ff}}{4},\frac{d_{ff}}{2},d_{ff}\}$.
With the nested MatFormer blocks $T_{1},T_{2}\dots T_{g}$, we can combine these to form a MatFormer model, with $g$ nested submodels $\mathcal{M}_{1}\subset\mathcal{M}_{2}\dots,\subset\mathcal{M}_{g}$ where $\mathcal{M}_{i}\leftarrow[T_{i}]^{\times l}$, i.e., $\mathcal{M}_{i}$ is formed by stacking $T_{i}$ for $l$ layers. The input and output embedding matrices are shared across the models.

### v2

@@ Abstract
Foundation models are applied in a broad spectrum of settings with different inference constraints, from massive multi-accelerator clusters to resource-constrained standalone mobile devices. However, the substantial costs associated with training these models often limit the number of unique model sizes that can be offered. Consequently, practitioners are compelled to select a model that may not be optimally aligned with their specific latency and cost requirements. We present MatFormer [^1] MatFormer stands for [Uncaptioned image] Matryoshka Transformer, reflecting the model’s inherent nested structure., a novel Transformer architecture designed to provide elastic inference across diverse deployment constraints. MatFormer achieves this by incorporating a nested Feed Forward Network (FFN) block structure within a standard Transformer model. During training, we optimize the parameters of multiple nested FFN blocks with varying sizes, enabling the extraction of hundreds of accurate smaller models without incurring additional computational costs. We empirically validate the efficacy of MatFormer across different model classes (decoders and encoders) and modalities (language and vision), demonstrating its potential for real-world deployment. We show that a 850M decoder-only MatFormer language model (MatLM) allows us to extract multiple smaller models spanning from 582M to 850M parameters, each exhibiting better validation loss and one-shot downstream evaluations than independently trained counterparts. Furthermore, we observe that smaller encoders extracted from a universal MatFormer-based ViT (MatViT) encoder preserve the metric-space structure for adaptive large-scale retrieval. Finally, we showcase that speculative decoding with the accurate and consistent submodels extracted from MatFormer can lead to significant reduction in inference latency. Project website.
@@ 1 Introduction
Figure 1: MatFormer introduces nested structure into the Transformer’s FFN block & trains all the submodels, enabling free extraction of hundreds of accurate submodels for elastic inference.
@@
Large Foundation models [49, 45, 17] are deployed in a variety of settings with different compute and accuracy demands like real-time response on mobile phones or on multi-cluster GPUs for web-scale batch serving. However, typical model families provide only a few independently trained models of different sizes. For example, the Llama-2 family provides models with 7B, 13B, 34B, and 70B parameters [59]. So practitioners are forced to choose a smaller (and typically less accurate) model than their latency/cost budget. Alternatively, one can use compression/pruning to fit a bigger model in a given compute budget [19, 36, 53], but that requires additional training.
@@
We introduce MatFormer, a natively elastic Transformer [61] architecture that overcomes this challenge. MatFormer allows for training one universal model which can be used to extract hundreds of smaller submodels without any additional training cost (Figure 1). MatFormer is a general architecture that can be applied to encoders and decoders, is domain agnostic, and is compatible with standard foundation model training pipelines.
@@
MatFormer follows the principle of matryoshka representation learning [34], to introduce nested substructure inside the standard Transformer block. Formally, MatFormer defines Transformer blocks $T_{i}$, such that, $T_{1}\subset T_{2}\subset\dots\subset T_{g}$, where $g$ is the number of nested transformer blocks, and $T_{i}\subset T_{i+1}$ relation indicates that the parameters of $T_{i}$ are contained in those of $T_{i+1}$. MatFormer can induce such sub-structure in both the attention and the feedforward network (FFN) blocks of the Transformer (see Figure 1). Consider, say, a FFN block that has $d_{\text{ff}}$ neurons in the hidden layer. Then, MatFormer induces matryoshka structure on these neurons, where $T_{i}$ contains the first $m_{i}$ neurons and $1\leq m_{1}<m_{2}\dots<m_{g}=d_{\text{ff}}$ represent the number of neurons for each granularity or sub-model. Intuitively, this implies that the first $m_{1}$ neurons are “most significant” neurons as they belong to all the blocks followed by the next $m_{2}-m_{1}$, and so on.
@@
In a departure from related work (Section 2), despite optimizing for only $g$ granularities, we are able to extract exponentially more submodels post-training. Using the trained MatFormer blocks $T_{1},\dots,T_{g}$ at each layer, one can form new models by Mix’n’Match (Section 3.3), i.e., by taking an arbitrary combination of these blocks across layers. For example, in the first layer, one can select $T_{g}$, the largest block, choose $T_{2}$ in the second layer, and so on, forming $g^{l}$ different models (where $l$ is the number of layers). Surprisingly, in multiple settings, and for various model sizes, we observe that the extracted models indeed are accurate, with accuracy scaling with the size of the extracted model.
@@
We train Matformer-based decoder-only Language Models (MatLM) up to 850M parameters and observe that: (a) MatLMs explicitly trained with $g$ exponentially spaced granularities outperform validation loss and one-shot downstream evals of respective $g$ baseline models trained independently from scratch, (b) our extracted models using Mix’n’Match lie on the accuracy-vs-parameters trade-off curve generated by the $g$ explicitly trained models, (c) through scaling experiments we observe that the loss vs compute law for different MatFormer models remains similar to vanilla Transformer models across different granularities and (d) the submodels extracted from MatLM have highly consistent behavior that is highly desirable for inference optimizations and deployment across scales.
@@
We further study MatFormer-based ViT models (MatViT) and make similar observations. For example, MatViT-L/16 improves the accuracy of the standard ViT-L/16 model on ImageNet-1K, and the extracted sub-models all match or even perform better than the independently trained baselines. Furthermore, we demonstrate that, due to high consistency, MatViT models can be used as “elastic encoders” for adaptive image retrieval. That is, the metric-space of an image encoded by the universal (i.e. largest) MatViT model is roughly preserved by the nested submodels. Hence, based on query complexity, system load, and various other considerations, we can use one of the extracted MatViT encoders at inference time for retrieval on a fixed corpus encoded by the universal model – providing over $40\%$ less compute overhead with $<0.5\%$ drop in accuracy.
@@
    We introduce MatFormer, which incorporates a nested sub-structure within the standard Transformer and optimizes all the $g$ granularities to produce a single, universal elastic model.
@@
    We introduce Mix’n’Match, a simple heuristic with no computation overhead that finds optimal submodels within a given parameter budget, outperforming more complex NAS methods. This yields hundreds of accurate and consistent submodels without any training cost (Section 3).
@@ 2 Related Work
Transformers [61] have become the unifying model architecture for foundation models [7] across modalities like language [8], vision [17] and audio [47]. While powerful, the standard Transformer block is not natively elastic in a way that enables large-scale adaptive and flexible deployment across various resource constraints. To cater to the plethora of deployment requirements, existing solutions include training a family of models of varying sizes [49, 2], post-hoc efficiency techniques like quantization [19], pruning [36], and distillation [53]. However, these solutions often are specific to the single constraint at hand, and require additional training for each new downstream usecase. This makes them far from being a truly elastic solution for adaptive deployment. Lastly, Transformer based LLMs are often sped-up during inference with techniques like speculative decoding [39, 12] – that benefits from the smaller draft & the larger verifier models having similar behavior – or early exiting [54] to enable real-time deployment.
Table 1: Comparison of MatFormer with comparable techniques across training and inference. We emphasize that in contrast with earlier work, MatFormer requires optimizing fewer models to obtain an exponential number of models at inference time without the need for post-training or NAS. Moreover, MatFormer subnetworks are nested, allowing for adaptive retrieval & colocation of models during inference. Here, $l$ is the number of layers in the model and $\exp(l)$ refers to exponential in $l$.
N(Models Optimized) N(Models Obtained) Nested? Model Selection Post-training Needed? Architecture Decoder Model?
MatFormer O(1) $exp(l)$ Mix’n’Match × Transformer
OFA [9] $exp(l)$ $exp(l)$ NAS × CNN ×
Slimmable Networks [66] O(1) O(1) - × CNN ×
HAT [63] $exp(l)$ $exp(l)$ NAS CNN enc-dec
Sorted Network [60] $exp(l)$ $exp(l)$ Both ×
DynaBERT [27] $O(1)$ $O(1)$ Transformer ×
Obtaining multiple smaller models from a single model has been explored in the past [66, 65, 9, 23, 10] with most work focusing on CNN encoders. In this work, we focus on Transformers in decoder-only language models and pretrained vision models. Specifically, OFA [9] trains a teacher CNN model, and uses distillation to finetune randomly sampled submodels (not nested) in the universal student CNN model. Moreover, OFA focuses on small scale models to be deployed on end devices. In contrast, MatFormer doesn’t require distillation, thereby using substantially less memory, and uses nested models. Using nested models allows us to host multiple models together without significantly increasing the model’s memory footprint, which is advantageous for scenarios where we want to route queries through different sub-networks. Slimmable networks [66] jointly optimizes and provide limited preset widths. Universal Slimmable network [65] extends this to sample from a continuous search space of submodels and optimizes them jointly. In contrast, MatFormer samples and only optimizes one of the preset granularities. HAT  [63] trains a universal network only to learn relative performance for different architectures. For deployment, the authors use NAS to find the optimal architecture and train it from scratch before serving. In contrast, MatFormer requires no additional training and accurate subnetworks can be obtained using Mix’n’Match (Section 3.3) yielding results as good as NAS without the additional complexity. DynaBERT [27] jointly trains a fixed set of submodels, doesn’t introduce any search strategy and discusses only using explicitly trained granularities as submodels. As a result of joint optimization of all granularities, DynaBERT yields fewer gradient updates and hence suboptimal performance compared to MatFormer while using same compute and memory (Section 4).
@@
Moreover, we emphasize that while most work in this area optimizes for an exponential number of models, we optimize a small number of subnetworks ($g=4$) to obtain an exponential number of models at inference time which leads to significantly better accuracy for large datasets. More recently, some of this work has been extended to Transformer encoders [11, 52] for extracting sub-models in both static or dynamic settings. But they either fail at extending further to decoder-only language models ([52]) or perform suboptimal to MatFormer (Section 4) owing to differences in their training methodology. While not in the weight space, matryoshka representation learning [34] & FlexiViT [5] showcase elasticity in output & input spaces respectively by smoothly spanning deployment constraints with minimal overhead. MatFormer, in contrast, builds upon these works by creating nested structure within the weight space instead to enable truly elastic and adaptive Transformer-based (decoder & encoder) models that span all the accuracy-vs-compute tradeoff (statically or dynamically) with minimal changes and training overhead (Figure 1). SortedNet [60] is a concurrent work with similar goals which optimizes many sampled submodels (akin to prior work) unlike MatFormer’s optimization of a few (typically 4) nested submodels. We also note Flextron [64], a recent work that builds upon MatFormer by extending the nested elasticity in both MLP and Attention Heads simultaneously, tail patches the Matformer style training, and includes a router to automatically route each token within the different granularities in each layer.
In Table 1, we summarize the differences between MatFormer and the related work discussed. Among these works, we identified two central ideas commonly employed: jointly training multiple submodels and sampling random submodels. We consider DynaBERT [27] and Once-for-All (OFA) [9] as the respective most relevant prior works and compare them to MatFormer in Section 4.
@@ 3.1 MatFormer Structure
MatFormer defines $g$ Transformer blocks $T_{i}$, such that, $T_{1}\subset T_{2}\subset\dots\subset T_{g}$ where $T_{i}\subset T_{i+1}$ indicates that the parameters of $T_{i}$ are contained in those of $T_{i+1}$. While it is possible to impose such a structure on any part of the Transformer, we select the FFN block to define our method and present a majority of our experiments, as the model size and computational cost of a Transformer is dominated (around $60\%$ for LLMs and ViTs) by the FFN block (see Appendix C, and Appendix F.2 for experiments applying MatFormer to the attention block of the Transformer). So, in this work, we focus on inducing the MatFormer’s nested sub-structure in the FFN block. We then stack individual blocks (for $l$ layers) to form $g$ nested models ($\mathcal{M}_{1\dots g})$ with shared parameters i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{i+1}$.
The Transformer FFN block has a single hidden layer with $d_{\text{ff}}$ neurons and both input and outputs in $\mathbb{R}^{d_{\text{model}}}$, and fixed FFN ratio $:=d_{\text{ff}}/d_{\text{model}}$ (typically $\geq 4$). MatFormer introduces the matryoshka nested structure with $g$ granularities on the hidden representation of the FFN block. Concretely, a nested sub-block of the Transformer, $T_{i}$ contains the first $m_{i}$ neurons of the FFN and $1\leq m_{1}<\dots<m_{g}=d_{\text{ff}}$ represent the number of neurons for each granularity or sub-model. So, depending on the chosen granularity the FFN operation of $T_{i}$ i.e., $T_{i}^{\text{FFN}}$ on an input $x\in\mathbb{R}^{d_{\text{model}}}$ is:
@@
where the weight matrices of FFN are $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{d_{\text{ff}}\times d_{\text{model}}}$ and bias terms are omitted for simplicity. $\mathbf{W}_{1}[0:k]$ denotes the submatrix with the first $k$ rows of $\mathbf{W}_{1}$. Finally, $\sigma$ is a non-linearity often set to GELU [24] or squared ReLU [56]. In this work, we chose the $g=4$ exponentially spaced granularities with FFN ratios of $\{0.5,1,2,4\}$ i.e., the nested hidden neurons are of the sizes $\{\frac{d_{ff}}{8},\frac{d_{ff}}{4},\frac{d_{ff}}{2},d_{ff}\}$. We get $g$ nested submodels $\mathcal{M}_{1}\subset\mathcal{M}_{2}\dots,\subset\mathcal{M}_{g}$ where $\mathcal{M}_{i}\leftarrow[T_{i}]^{\times l}$, i.e., $\mathcal{M}_{i}$ is formed by stacking $T_{i}$ for $l$ layers. The input and output embedding matrices are shared across the models.
We note that we can form a similar sub-structure on the attention heads, with the heads being organized from “most” to “least” significant, where the more significant heads are shared by more sub-models. That is, we use the first $m_{i}$ attention heads for the $i$th granularity. We can also introduce this sub-structure in the token embedding ($d_{\text{model}}$) supplied to each Transformer block.